# Coupled Axial and Transverse Currents Method for Finite Element Modelling of Periodic Superconductors

### v2

@@ Abstract
In this paper, we propose the Coupled Axial and Transverse currents (I) (CATI) method, as an efficient and accurate finite element approach for modelling the electric and magnetic behavior of periodic composite superconducting conductors. The method consists of a pair of two-dimensional models coupled via circuit equations to account for the conductor geometrical periodicity. This allows to capture three-dimensional effects with two-dimensional models and leads to a significant reduction in computational time compared to conventional three-dimensional models. After presenting the method in detail, we verify it by comparison with reference finite element models, focussing on its application to twisted multifilamentary superconducting strands. In particular, we show that the CATI method captures the transition from uncoupled to coupled filaments, with accurate calculation of the interfilament coupling time constant. We then illustrate the capabilities of the method by generating a detailed loss map and magnetization curves of given strand types for a range of external transverse magnetic field excitations, with and without transport current.
  • May 1, 2024
@@
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
[…]
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Reduced order models based on a helicoidal change of variables [19, 20, 21], leading to special two-dimensional (2D) models, are one possibility, but they are limited to helicoidally symmetric conductor cross-sections and remain challenging to implement in the case of transverse field excitation, i.e., perpendicular to the strand axis, with nonlinear materials [21]. Treating such excitations is however crucial in the context of AC loss calculation.
@@
The method proposed in [25, 26, 27] reproduces the effect of coupling currents in the conducting matrix by introducing equivalent resistances between the filaments, accounting for the twist pitch length of the wire. This allows to reproduce coupled filament regimes without using a 3D model, but a simple, yet appropriate, 2D approach. The method presented in this contribution is inspired by that work.
We propose to consider a pair of 2D models: one for calculating axial currents, flowing along the wire axis, and one for modelling coupling currents, or transverse currents, flowing perpendicular to the wire axis, in the matrix in-between the superconducting filaments. These models are then coupled via circuit equations that represent the effect of the periodic geometry of the conductor. We refer to this method as CATI, standing for Coupled Axial and Transverse currents (I).
@@
Another advantage of the method is the relative ease of its implementation. It can be implemented in any FE software allowing for field-circuit coupling. In this work, the method is implemented in GetDP [28] within FiQuS [29], developed at CERN as part of the STEAM framework [30]. Geometry and mesh generation is performed by Gmsh [31]. All the software is open-source and free to use, and the main model files are (will be) available online111[Online] https://cern.ch/fiqus..
In Section 2, we present the CATI method and describe in detail its practical implementation. We also highlight its underlying assumptions and limitations. We then verify and assess the range of applicability of the approach in two steps. First, in Section 3, we consider a linear problem with constant material properties, and discuss the accuracy of the method, focussing on the coupling current dynamics. Second, in Section 4, we consider a nonlinear problem with realistic field-dependent material parameters and verify the prediction of the CATI method against those of a 3D FE reference model. Finally, in Section 5, we exploit the method and demonstrate its capabilities on a helicoidally symmetric strand and on a non-helicoidally symmetric, but periodic, wire-in-channel geometry.
@@ 2 CATI Method
We consider a 2D cross-section of a periodic composite conductor, perpendicular to the conductor main axis $\hat{\boldsymbol{e}}_{z}$. The cross-section contains $N_{\text{f}}\in\mathbb{N}_{0}$ filaments $\Omega_{\text{f},i}$, with $i\in F=\{1,\dots,N_{\text{f}}\}$. The filament positions are assumed to be periodically appearing in exactly the same locations after a given displacement along $\hat{\boldsymbol{e}}_{z}$. For a round strand such as the one shown in Figs. 1 and 2, this is achieved after twisting the strand with a twist pitch length $p$.
Figure 1: Computational domain for the cross-section of a superconducting strand ($N_{\text{f}}=54$). The 3D geometry is generated by a rotated extrusion of the cross-section along $\hat{\boldsymbol{e}}_{z}$ with twist pitch length $p$ (see Fig. 2). The curved arrow represents the rotation of filament $i$ to filament $\mathcal{S}(i)$ along the periodicity length $\ell=p/6$.
Figure 2: 3D representation of the 54-filament composite strand whose transverse cross-section is represented in Fig. 1. The cylindrical outline represents the boundary of the conducting matrix in which the filaments are embedded.
@@
The CATI method is based on the assumption that there is a periodicity of the cross-sections along $\hat{\boldsymbol{e}}_{z}$. Let $\ell$ be the periodicity length, that is, the distance along $\hat{\boldsymbol{e}}_{z}$ after which every filament takes the position of another one. Let the link between the positions in successive cross-sections, separated by a distance $\ell$, be described by a permutation operator $\mathcal{S}$, defining a bijection of the set of filaments onto itself,
@@
For example, with the geometry in Figs. 1 and 2, the smallest possible value of $\ell$ is $\ell=p/6$, and filament $\Omega_{\text{f},\mathcal{S}(i)}$ is the image of filament $\Omega_{\text{f},i}$ under a $62\pi/$ 6 rotation around the strand center (see the dashed arrow in Fig. 1).
The periodicity does not need to be created by twisting the geometry; the method only requires a periodic structure of the cross-sections. Examples of such periodic, but not twisted, cross-sections include common cable types, like Rutherford [32] or Roebel [33]. The general concept of the method equally applies to them. However, they are outside of the scope of this contribution.
@@ 2.1 CATI method equations
The CATI method consists of a pair of 2D models: (i) a model describing axial currents (along $\hat{\boldsymbol{e}}_{z}$) in $\Omega_{\text{c}}=\Omega_{\text{f}}\cup\Omega_{\text{m}}$, and the associated transverse magnetic field (perpendicular to $\hat{\boldsymbol{e}}_{z}$) in $\Omega=\Omega_{\text{c}}\cup\Omega_{\text{c}}^{\text{C}}$, and (ii) a model describing transverse currents (perpendicular to $\hat{\boldsymbol{e}}_{z}$) in $\Omega_{\text{m}}$ only. This is illustrated in Fig. 3. In the following, we refer to these two models as AI and TI, standing for axial currents and transverse currents, respectively. Net currents and voltages are introduced for each model, they are referred to as global quantities. The AI and TI models are coupled via these global quantities using circuit equations. In this section, we define the weak form of the problem in the continuous setting.
@@ 2.1.1 Axial currents formulation
We define $\boldsymbol{j}$ and $\boldsymbol{e}$ as axial vectors, and $\boldsymbol{h}$ and $\boldsymbol{b}$ as transverse vectors. The chosen weak form of the model is that of a classical 2D $h$-$\phi$-formulation, in order to ensure the numerical efficiency in the treatment of the superconductor power law for the resistivity [35]. The 2D formulation is integrated along $\hat{\boldsymbol{e}}_{z}$ over a length $2\ell$, assuming that the solution is constant over this length. The choice of integrating over $2\ell$ instead of $\ell$ is discussed in Section 2.1.3.
@@ 2.1.3 Circuit-coupling equations
The AI and TI models involve $24N_{\text{f}}+$ 2 global quantities. There are already $12N_{\text{f}}+$ 1 equations linking them in Eqn. (2.1.1) and (9), so that only $12N_{\text{f}}+$ 1 equations are still to be defined. One equation consists in setting directly either the transport current $I_{\text{t}}$ or the associated voltage $V_{\text{t}}$. The remaining $2N_{\text{f}}$ equations are defined below; they couple the AI and TI models and are the core the CATI method. They rely on the periodicity of the geometry of the conductor.
@@
We illustrate the meaning of Eqn. (10) and (11) on a simple example with $N_{\text{f}}=6$ filaments in Fig. 4. Using the periodicity of the cross-sections, the difference between the currents $I_{3}$ and $I_{1}$ is equal to the accumulated current $\tilde{I}_{2}$ over the length $2\ell$. Similarly, the difference between the voltages $\tilde{V}_{3}$ and $\tilde{V}_{1}$ is equal to the accumulated voltage $V_{2}$ over the length $2\ell$. The AI and TI models can therefore be interpreted as being solved on staggered cross-sections. By periodicity, all these cross-sections are equivalent.
@@
Note that it is possible to reduce the integration along $z$ from $2\ell$ down to $\ell$, so that the AI and TI models are solved on distinct cross-sections. In the case of Fig. 4, this leads to TI cross-sections that are rotated by an angle of $\pi/6$ with respect to the AI ones. For the geometries considered in this contribution, we observed that the obtained results (with adapted corrected length, as discussed in Section 2.2) are almost identical.
@@ 2.1.4 Underlying assumptions
The CATI method requires a periodic geometry of the conductor, whose cross-section repeats itself after a given length in the main direction of the conductor $\hat{\boldsymbol{e}}_{z}$. Along this length, filaments must move from one location to another, and hence be tilted with respect to $\hat{\boldsymbol{e}}_{z}$. Because purely 2D models are used to describe the fields, one main assumption of the CATI method is that the effect of the tilt angle on the field distribution can be neglected.
@@
The CATI method also assumes that the filament resistivity is sufficiently small compared to the matrix resistivity for considering that the TI model voltage $v$ is constant on each filament boundary. However, if this is not the case, axial currents will be more evenly shared among the matrix and the filaments, and the contribution of coupling currents to total losses is therefore expected to decrease. In the extreme case of a quench, when the filaments become relatively highly resistive, most of the current flows in the matrix, and coupling currents become negligible. These situations can be handled by the CATI method, as it allows for axial current flow in the matrix.
The heart of the method lies in the circuit-coupling equations. They describe the evolution of fields and global quantities in the filaments along $\hat{\boldsymbol{e}}_{z}$. This evolution is represented by the global quantities of distinct filaments within a single cross-section, exploiting the periodicity of it. For the example in Fig. 4, the evolution of global quantities of filaments can be decomposed in up to $6$ steps before a complete rotation is performed. A minimum of 2 steps is necessary for the method to be relevant.
@@ 2.2 Length correction factor
The use of 2D models leads to assuming that the filaments carry constant currents $I_{i}$ and maintain constant voltages $\tilde{V}_{i}$ over a distance $2\ell$ along the wire, in the AI and TI models, respectively, after which the position of one filament suddenly changes to the next one, and so are the values of the currents and voltages. Similarly, the voltages $V_{i}$ and the currents $\tilde{I}_{i}$ are obtained by simple multiplication of 2D solutions by $2\ell$.
@@
To account for this effect, a correction factor can be evaluated analytically. With $\alpha=2\pi/p$, at a given constant position $\bar{z}$, we have
@@
where integrands are symmetrically integrated around $\bar{z}$ due to the symmetry of the circuit coupling equations. Note that Eq. (2.2) does not depend on $\bar{z}$. This implies that the overestimation of the twist effect is identical for all filaments. For most conductors used in practice, $\ell=p/6$, and the correction factor is equal to $\approx 0.8270$.
@@
We will show in Section 3 that this correction strongly improves the accuracy of the model, helping to better reproduce the analytical time constant associated with the coupling currents.
In the alternative case where the equations are integrated over length $\ell$ instead of $2\ell$, leading to the AI and TI models to be solved on rotated distinct cross-sections, $2\ell$ must be replaced by $\ell$ in Eq. (2.2). In the common case in which $\ell=p/6$, this leads to a correction factor of $\approx 0.9549$.
A similar approach can be followed for other conductor types, for which quantities are expected to follow a non-sinusoidal evolution along the conductor. It is however not guaranteed in such cases that a common correction factor can be derived for the whole cross-section, as is the case above with the invariance of Eq. (2.2) with respect to $\bar{z}$. An average correction factor may be considered in such cases.
@@ 2.3 Coupling current dynamics
with $\mathcal{H}_{\perp}(\Omega_{\text{m}})$ the subspace of $H(\textrm{{curl}};\Omega_{\text{m}})$ containing axial vector fields (parallel to $\hat{\boldsymbol{e}}_{z}$). In practice, we gauge the function space by imposing a zero average value of $\boldsymbol{h}_{\text{s}}$ over $\Omega_{\text{m}}$, via a Lagrange multiplier [38]. Note that, locally, $\textrm{{curl}}\,\boldsymbol{h}_{\text{s}}\neq-\sigma\,\textrm{{grad}}\,v$. Indeed, as mentioned above, an axial field alone cannot reproduce an transverse coupling current distribution with net current exchanges between the filaments. The field $\boldsymbol{h}_{\text{s}}$ is only an approximation of $\boldsymbol{h}_{\text{cc}}$ that belongs to $\mathcal{H}_{\perp}(\Omega_{\text{m}})$.

### v3

@@ Abstract
In this paper, we propose the Coupled Axial and Transverse currents (I) (CATI) method, as an efficient and accurate finite element approach for modelling the electric and magnetic behavior of periodic composite superconducting conductors. The method consists of a pair of two-dimensional models coupled via circuit equations to account for the conductor geometrical periodicity. This allows to capture three-dimensional effects with two-dimensional models and leads to a significant reduction in computational time compared to conventional three-dimensional models. After presenting the method in detail, we verify it by comparison with reference finite element models, focussing on its application to twisted multifilamentary superconducting strands. In particular, we show that the CATI method captures the transition from uncoupled to coupled filaments, with accurate calculation of the interfilament coupling time constant. We then illustrate the capabilities of the method by generating detailed loss maps and magnetization curves of given strand types for a range of external transverse magnetic field excitations, with and without transport current.
  • June 24, 2024
@@ 1 Introduction
Reduced order models based on a helicoidal change of variables [19, 20, 21], leading to special two-dimensional (2D) models, are one possibility, but they are limited to helicoidally symmetric conductor cross sections and remain challenging to implement in the case of transverse field excitation, i.e., perpendicular to the strand axis, with nonlinear materials [21]. Treating such excitations is however crucial in the context of AC loss calculation.
@@
The method proposed by T. Satiramatekul and F. Bouillault in [25, 26, 27] reproduces the effect of coupling currents in the conducting matrix by introducing equivalent resistances between the filaments, accounting for the twist pitch length of the wire. This allows to reproduce uncoupled, partially coupled, and coupled filament regimes without using a 3D model, but a simple, yet appropriate, 2D approach. The method presented in this contribution is inspired by that work.
We propose to consider a pair of 2D models: one for calculating axial currents, flowing along the wire axis, and one for modelling coupling currents, or transverse currents, flowing perpendicular to the wire axis, in the matrix in-between the superconducting filaments. We then couple these models using circuit equations that represent the effect of the periodic geometry of the conductor. We refer to this approach as the CATI method, standing for Coupled Axial and Transverse currents (I) method.
@@
Another advantage of the method is the relative ease of its implementation. It can be implemented in any FE software allowing for field-circuit coupling. In this work, the method is implemented in GetDP [28] within FiQuS [29], developed at CERN as part of the STEAM framework [30]. Geometry and mesh generation is performed by Gmsh [31]. All the software is open-source and free to use111[Online] https://cern.ch/fiqus.. The input files for the simulations are provided in [32] so that the CATI results can be reproduced.
In Section 2, we present the CATI method, describe in detail its practical implementation, and highlight its underlying assumptions and limitations. We then verify and assess the range of applicability of the approach in two steps. First, in Section 3, we consider a linear problem with constant material properties, and discuss the accuracy of the method, focussing on the coupling current dynamics. Second, in Section 4, we consider a nonlinear problem with realistic field-dependent material parameters and verify the prediction of the CATI method against those of a 3D FE reference model. Finally, in Section 5, we exploit the method and demonstrate its capabilities on a helicoidally symmetric strand and on a non-helicoidally symmetric, but periodic, wire-in-channel geometry.
@@ 2 CATI Method
We consider a 2D cross section of a periodic composite conductor, perpendicular to the conductor main axis $\hat{\boldsymbol{e}}_{z}$. The cross section contains $N_{\text{f}}\in\mathbb{N}_{0}$ filaments $\Omega_{\text{f},i}$, with $i\in F=\{1,\dots,N_{\text{f}}\}$. The filament positions are assumed to be periodically appearing in exactly the same locations after a given displacement along $\hat{\boldsymbol{e}}_{z}$. For a round strand such as the one shown in Figs. 1 and 2, this is achieved after twisting the strand with a twist pitch length $p$.
Figure 1: Computational domain for the cross section of a superconducting strand ($N_{\text{f}}=54$). The 3D geometry is generated by a rotated extrusion of the cross section along $\hat{\boldsymbol{e}}_{z}$ with twist pitch length $p$ (see Fig. 2). The curved arrow represents the rotation of filament $i$ to filament $\mathcal{S}(i)$ along the periodicity length $\ell=p/6$.
Figure 2: 3D representation of the 54-filament composite strand whose transverse cross section is represented in Fig. 1. The cylindrical outline represents the boundary of the conducting matrix in which the filaments are embedded.
@@
The CATI method is based on the assumption that there is a periodicity of the cross sections along $\hat{\boldsymbol{e}}_{z}$. Let $\ell$ be the periodicity length, that is, the distance along $\hat{\boldsymbol{e}}_{z}$ after which every filament takes the position of another one. Let the link between the positions in successive cross sections, separated by a distance $\ell$, be described by a permutation operator $\mathcal{S}$, defining a bijection of the set of filaments onto itself,
@@
For example, with the geometry in Figs. 1 and 2, the smallest possible value of $\ell$ is $p/6$, and filament $\Omega_{\text{f},\mathcal{S}(i)}$ is the image of filament $\Omega_{\text{f},i}$ under a $62\pi/$ 6 rotation around the strand center (see the dashed arrow in Fig. 1).
The periodicity does not need to be created by twisting the geometry; the method only requires a periodic structure of the cross sections. Examples of such periodic, but not twisted, cross sections include common cable types, like Rutherford [33] or Roebel [34]. The general concept of the method applies to them. However, they are outside of the scope of this contribution.
@@
In Section 2.1, we define the strong and weak forms of the method and discuss the assumptions that are made. In Sections 2.2 and 2.3 we propose two improvements of the method. Both are optional in the implementation, but both improve significantly the accuracy of the model. Finally, in Section 2.4, we focus on the practical implementation of the method, by describing in detail the spatial discretization of the unknown fields with finite elements.
@@ 2.1 CATI method equations
The CATI method consists of a pair of 2D models: (i) a model describing axial currents (along $\hat{\boldsymbol{e}}_{z}$) in $\Omega_{\text{c}}=\Omega_{\text{f}}\cup\Omega_{\text{m}}$, and the associated transverse magnetic field (perpendicular to $\hat{\boldsymbol{e}}_{z}$) in $\Omega=\Omega_{\text{c}}\cup\Omega_{\text{c}}^{\text{C}}$, and (ii) a model describing transverse currents (perpendicular to $\hat{\boldsymbol{e}}_{z}$) in $\Omega_{\text{m}}$ only. This is illustrated in Fig. 3. In the following, we refer to them as the AI and TI models, standing for axial currents and transverse currents models, respectively. Net currents and voltages are defined for each filament in both models, they are referred to as global quantities. The AI and TI models are coupled via these global quantities using circuit equations that are defined in order to account for the periodicity of the geometry, hence of its three-dimensional aspect. In particular, the periodicity length $\ell$ is explicitly present in the governing equations.
In this section, we define the strong and weak forms of the problem in the continuous setting. We successively introduce the AI model, the TI model, and the circuit equations. We conclude by discussing the assumptions behind the equations.
@@ 2.1.1 Axial currents formulation
We define $\boldsymbol{j}$ and $\boldsymbol{e}$ as axial vectors, and $\boldsymbol{h}$ and $\boldsymbol{b}$ as transverse vectors. The chosen weak form of the model is that of a classical 2D $h$-$\phi$-formulation, in order to ensure the numerical efficiency in the treatment of the superconductor power law for the resistivity [36]. The 2D formulation is integrated along $\hat{\boldsymbol{e}}_{z}$ over a length $2\ell$, assuming that the solution is constant over this length.
@@
In reality, the solution is not constant over the distance $2\ell$, but rather varies smoothly along the wire. We discuss how to correct for this assumption in Section 2.2. Also, integrating over $2\ell$ is a choice. Integrating over $\ell$ instead leads to an equally valid model and is discussed in Section 2.1.3.
@@ 2.1.3 Circuit-coupling equations
The AI and TI models involve $24N_{\text{f}}+$ 2 global quantities. There are already $12N_{\text{f}}+$ 1 equations linking them in Eqn. (2.1.1) and (9), so that only $12N_{\text{f}}+$ 1 equations are still to be defined. One equation consists in setting directly either the transport current $I_{\text{t}}$ or the associated voltage $V_{\text{t}}$. The remaining $2N_{\text{f}}$ equations are defined below; they couple the AI and TI models and are the core of the CATI method. They rely on the periodicity of the geometry of the conductor.
@@
We illustrate the meaning of Eqn. (10) and (11) on a simple example with $N_{\text{f}}=6$ filaments in Fig. 4. Using the periodicity of the cross sections, the difference between the currents $I_{3}$ and $I_{1}$ is equal to the current $\tilde{I}_{2}$, accumulated over the length $2\ell$. Similarly, the difference between the voltages $\tilde{V}_{3}$ and $\tilde{V}_{1}$ is equal to the voltage $V_{2}$, accumulated over the length $2\ell$. The AI and TI models can therefore be interpreted as being solved on staggered cross sections. By periodicity, all these cross sections are equivalent, which allows us to consider only one of them.
@@
Note that it is possible to reduce the integration along $z$ from $2\ell$ down to $\ell$, so that the AI and TI models are solved on distinct cross sections. In the case of Fig. 4, this leads to TI cross sections that are rotated by an angle of $\pi/6$ with respect to the AI ones. For the geometries considered in this contribution, we observed that the obtained results (with adapted corrected length, as discussed in Section 2.2) are almost identical.
@@ 2.1.4 Underlying assumptions
The CATI method requires a periodic geometry of the conductor, whose cross section repeats itself after a given length in the main direction of the conductor $\hat{\boldsymbol{e}}_{z}$. Along this length, filaments must move from one location to another, and hence be tilted with respect to $\hat{\boldsymbol{e}}_{z}$. Because purely 2D models are used to describe the fields, one main assumption of the CATI method is that the effect of the tilt angle on the field distribution can be neglected.
@@
The CATI method also assumes that the filament resistivity is sufficiently small compared to the matrix resistivity for considering that the TI model voltage $v$ is constant on each filament boundary. However, if this is not the case, axial currents will be more evenly shared among the matrix and the filaments, and the contribution of coupling currents to total losses is therefore expected to decrease. In the extreme case of a quench, when the filaments become relatively highly resistive, most of the current flows in the matrix, and coupling currents become negligible. These situations can be handled by the CATI method, as it describes the axial current flow in the matrix.
The heart of the method lies in the circuit-coupling equations. They describe the evolution of fields and global quantities in the filaments along $\hat{\boldsymbol{e}}_{z}$. This evolution is represented by the global quantities of distinct filaments within a single cross section, exploiting the periodicity of it. For the example in Fig. 4, the evolution of global quantities of filaments can be decomposed in up to $6$ steps before a complete rotation is performed. A minimum of 2 steps is necessary for the method to be relevant.
@@ 2.2 Length correction factor
The use of 2D models leads to assuming that the filaments carry constant currents $I_{i}$ and maintain constant voltages $\tilde{V}_{i}$ over a distance $2\ell$ along the wire, in the AI and TI models, respectively, after which the position of one filament suddenly changes to the next one, and so do the values of the currents and voltages. Similarly, the voltages $V_{i}$ and the currents $\tilde{I}_{i}$ are obtained by simple multiplication of 2D solutions by $2\ell$.
@@
To account for this effect, a correction factor can be evaluated analytically. With $\alpha=2\pi/p$, at a given position $\bar{z}$, we have
@@
where integrands are symmetrically integrated around $\bar{z}$ due to the symmetry of the circuit coupling equations. The fact that Eq. (2.2) does not depend on $\bar{z}$ is convenient, as this implies that the overestimation of the twist effect is identical for all filaments, and can then in principle be corrected by a single correction factor. For many conductors used in practice, the filament structure is hexagonal, so that $\ell=p/6$ and the correction factor is equal to $\approx 0.8270$.
@@
The effect of the modification is to directly rescale the global quantities. We will show in Section 3 that this correction strongly improves the accuracy of the model, helping to better reproduce the analytical time constant associated with the coupling currents.
In the alternative case where the equations are integrated over length $\ell$ instead of $2\ell$, leading to the AI and TI models to be solved on rotated distinct cross sections, $2\ell$ must be replaced by $\ell$ in Eq. (2.2). In the common case in which $\ell=p/6$, this leads to a correction factor of $\approx 0.9549$.
A similar approach can be followed for other conductor types, for which quantities are expected to follow a non-sinusoidal evolution along the conductor. It is however not guaranteed in such cases that a common correction factor can be derived for the whole cross section, as is the case above with the invariance of Eq. (2.2) with respect to $\bar{z}$. An average correction factor may be considered in such cases.
@@ 2.3 Coupling current dynamics
with $\mathcal{H}_{\perp}(\Omega_{\text{m}})$ the subspace of $H(\textrm{{curl}};\Omega_{\text{m}})$ containing axial vector fields (parallel to $\hat{\boldsymbol{e}}_{z}$). In practice, we gauge the function space by imposing a zero average value of $\boldsymbol{h}_{\text{s}}$ over $\Omega_{\text{m}}$, via a Lagrange multiplier [39]. Note that, locally, $\textrm{{curl}}\,\boldsymbol{h}_{\text{s}}\neq-\sigma\,\textrm{{grad}}\,v$. Indeed, as mentioned above, an axial field alone cannot reproduce a transverse coupling current distribution with net current exchanges between the filaments. The field $\boldsymbol{h}_{\text{s}}$ is only an approximation of $\boldsymbol{h}_{\text{cc}}$ that belongs to $\mathcal{H}_{\perp}(\Omega_{\text{m}})$.